OpenStudy (anonymous):

Solve the equation 2cos^2x = sin2x , giving your answer in terms of π. π

OpenStudy (anonymous):

sorry it took trillion years in page loading, you know OS is damn slow.... \[2\cos^2x=2sinx.cosx => \cos x=\sin x=>tanx=1, x=\frac{ 3 \pi }{ 4 }\]